# Who’s in and who’s out? A case study of multimodal CLIP-filtering in DataComp

### v1

@@ 1. Introduction
Text and image datasets in machine learning have grown to the scale of billions (Devlin et al., 2018; Radford et al., 2019; Schuhmann et al., 2022) in order to build larger language models and text-to-image generative models (Hoffmann et al., 2022; AI, 2024). To reach such magnitudes, datasets are being increasingly scraped from the web (Raffel et al., 2020; Gadre et al., 2024), through archives such as Common Crawl (Crawl, 2024). While data from these web dumps are expansive, they are not expertly curated and may reduce the efficacy of downstream models. Moreover, it is infeasible to manually assess every data point, which therefore requires at-scale methods to remove data that curators deem undesirable.
To address the issues introduced by these large-scale data gathering practices, machine learning practitioners and researchers have increasingly begun to rely on automated data filtering to improve training efficiency and discard what they refer to as “low-quality” data. Specifically, for multimodal data, researchers have applied the OpenAI CLIP model to filter image-text data obtained from Common Crawl. As depicted in Figure 1, this CLIP-filtering method (referring to the use of the OpenAI CLIP model (Radford et al., 2021)) assesses similarity between an image and its corresponding text within the model’s embedding space. This filtering technique was used to create the LAION-400M (Schuhmann et al., 2021), LAION-5B (Schuhmann et al., 2022), and DataComp-1B (Gadre et al., 2024) datasets with sizes of 400 million, 5 billion, and 1 billion respectively, which were then used to train other open-source CLIP models. Because these models obtained state-of-the-art performance on several zero-shot image classification tasks, researchers have concluded CLIP-filtering to be the “most performant” method, as opposed to other filtering techniques that use caption length or image size (Gadre et al., 2024; Fang et al., 2023).
@@
At the same time, it is unclear what role CLIP-filtering plays on downstream text-to-image models, especially given the demographic biases embedded in OpenAI CLIP itself (Agarwal et al., 2021). While there have been speculations on the demographic biases of CLIP-filtering (Birhane et al., 2021), there has not been any exploration to assess exactly how the filter changes the demographic makeup of the training dataset. As such, we decide to examine this initial stage of dataset curation because of its potential impact on downstream models.
@@ 1.1. Contributions
We break down our investigation to address the following three questions: First: Which demographic groups are disproportionately excluded by CLIP-filtering? Different demographic groups being filtered at different rates implies a form of representational harm (Shelby et al., 2023). A filter that systematically omits data from a certain group may therefore worsen the performance of the resulting model on that group. As a result, we aim to quantify this breakdown between unfiltered and filtered data across various demographic dimensions. The need to assess sociodemographic attributes across multiple modalities motivates our next question: How can we measure sociodemographic attributes at scale within an image-text dataset? We examine a broad range of factors like gender, age, race, religion, sexuality, language, or geographic region. While by no means comprehensive, we infer attributes that appear in the text and image components that can easily be related to people. In addition to examining filtering discrepancies by demographics, we also examine discrepancies by data source: What types of websites are considered “high-quality” by the CLIP filter and what are the implications of certain websites being included? To the extent that it is valuable to know whether datasets are globally representative, we ask whether data from certain geographic regions or time periods of the internet are treated differently by CLIP-filtering, as well as which websites have data that are more likely to pass through the filter.
@@
    Filter discrepancies: Across multiple modalities, different imputed sociodemographic groups are filtered at different rates. For example, CLIP-filtering is more likely to exclude data relating to LGBTQ+ identity and non-Western regions.
@@
    Evaluation tools: To address our research questions, we also develop novel methodologies to effectively audit an image-text dataset at scale — focusing specifically on the multimodality of this data. We make our analysis code available at https://github.com/hongrachel/clip-filtering-bias/.
Overall, this existing data filtering method is a clear example of a value-embedded machine learning practice (Birhane et al., 2022). From our results, we see that CLIP-filtering does in fact substantially change the makeup of the final dataset. This demonstrates more broadly the idea from Gururangan et al. (2022) that data filtering must encode a specific ideology of what constitutes as “high-quality data.” Prior works on image-text filtering assume that “quality” is often intrinsic to the data and that this “noise” to be weeded out is vaguely defined (Wenzek et al., 2019). The LAION and DataComp authors demonstrate the efficacy of the CLIP filter by selecting specific downstream metrics for specific datasets such as zero-shot ImageNet classification accuracy, rather than the societal impacts of downstream models (Schuhmann et al., 2021; Gadre et al., 2024). Here, we investigate a case where filtering is designed around particular assumptions of what data quality and performance should be without additional justification. Despite the supposed objectivity of these decisions, our findings prove that filtering embeds societal norms as to what should and should not be excluded. Following our analysis, we further elaborate on these implications and give recommendations for designing data filters in the future.
@@ 3. Related Work
We highlight previous work on CLIP evaluations and data filtering methods that are relevant to our study. We aim to contribute to the existing discussion on data filtering bias through a novel case study of multimodal filtering.
@@ 3.2. CLIP-filtered dataset audits
Some recent works have audited the LAION datasets in order to understand the nature of the content within these large-scale datasets. Recently, Thiel (2023) uses hash-based detection and other computational techniques to identify thousands of CSAM (Child Sexual Abuse Material) images in LAION-5B, which resulted in the takedown of the LAION-5B dataset in December, 2023 (Verma and Harwell, 2023). In addition, Birhane et al. (2021, 2023) have found significant rates of hateful content, explicit images, stereotypes, and racial slurs in both the LAION-400M and LAION-2B-en datasets. Given the biases of CLIP, they illustrate hypothetical examples in which misogynist or racist descriptions would pass the CLIP filter but reasonable, benign descriptions would be excluded (Birhane et al., 2021). The question then remains: What, exactly, does CLIP-filtering include and exclude? Our work closely follows these prior works to study the problematic content in large-scale image-text datasets that pass through the filtering step. We aim to extend their findings to understand the impact of filtering on demographic representation at scale.
@@ 3.3. Data filtering bias
There also has been much discussion on the systematic exclusion of data from marginalized identities in the data cleaning stage more broadly. Bender et al. (2021) caution early on that filtering can suppress discourse from marginalized identities and compound power imbalances in text data collection practices. From a sociotechnical lens, Muller and Strohmayer (2022) describe how data cleaning can be a method of erasure that is difficult to detect. Desai et al. (2023) draw on an archival perspective to examine the social value judgements embedded in filtering choices.
We are inspired by the approaches in these prior filtering evaluations and aim to apply similar methods to multimodal data. To our knowledge, there has been no work assessing data filtering bias at scale in the context of image-text datasets.
@@ 3.5. Downstream text-to-image model harms
There has been ample work that evaluate the societal issues of downstream models trained on CLIP-filtered data. Luccioni et al. (2024), for instance, find that Stable Diffusion (which trains on LAION-5B obtained via CLIP-filtering (Rombach et al., 2022)), underrepresents marginalized identities in their image generations. Moreover, researchers have demonstrated that Stable Diffusion generates unsafe, hateful, or sexualized content at significantly high rates (Qu et al., 2023; Wolfe et al., 2023). Exposure to stereotypical and problematic imagery in general has been shown to shape people’s beliefs and behavior (Branley and Covey, 2017; Slusher and Anderson, 1987), and this becomes more pressing as these models grow increasingly popular. These recent findings thus all illustrate the incredibly harmful consequences of downstream models, so it is important to understand where these issues may arise in the ML pipeline.
@@ 4. Approach
In this section, we describe how we audit the CLIP-filtering step in DataComp and obtain demographic annotations and website groupings. We release our code at https://github.com/hongrachel/clip-filtering-bias/.
4.1. CLIP filter
We replicate the CLIP-filtering step for the small version of CommonPool of 12.8 million image-text pairs. We use the CLIP ViT-L/14 model and set the threshold as $0.243$ to select the top $30\%$ of pairs according to CLIP similarity score as was done in prior work (Schuhmann et al., 2022; Gadre et al., 2024). This results in a filtered dataset with 3.84 million image-text pairs.
@@ 4.2. Imputed demographic group annotations
In this manner, we infer demographic group or geographic region across multiple modalities in the CommonPool dataset: from the text content, the image content, and the source URL of the sample. As described in Section 5, this includes methods such as searching for mentions of keywords relating to demographic identity or applying an external face attribution predictor. We emphasize these findings do not represent behavior of this filtering according to true demographics; instead, they reveal how filtering interacts with imputed demographic attributes. We elaborate on the limitations of imputed demographics in Section 4.4. Findings here still have broad implications, as imputed demographics can correlate, although imperfectly, with socially salient populations (Elliott et al., 2009).
@@ 4.4. Limitations
We recognize numerous limitations to the sociodemographic imputation and source analysis techniques that we use. Third-party gender predictors, for instance, have been demonstrated to have disparate error rates across sensitive attributes and can rely on biases present in the face detection step (Buolamwini and Gebru, 2018; Schwemmer et al., 2020). Existing vision datasets have also been shown to have inconsistent definitions of race (Khan and Fu, 2021), and many gender annotation methods only give binary labels as “Male” and “Female” which conflates gender with sex and ignores non-binary individuals (Valdes, 1996). We acknowledge that gender and race are not inherently visual components, but rather socioculturally defined (Butler, 2013; Lopez, 1995). Identity keywords in the text also are rough proxies for the presence of sociodemographic attributes in the image or text.
@@ 5. Demographic Group Analysis
This section describes the demographic group imputation methods along both text and image modalities and presents their corresponding results.
@@ 5.1. Identity keyword
(a) Pass rate by identity keyword
@@
These differences in pass rates for certain intersections may relate to existing societal stereotypes. For instance, Latina, Asian American, and Pacific Islander women have been overtly sexualized and objectified in culture and media throughout US history (Azhar et al., 2021; Shimizu, 2007). Higher pass rates reinforce this notion, where data with text relating to Latina and Asian women may be overrepresented in the filtered dataset.
@@
The aforementioned gender stereotype associations give further evidence that CLIP-filtering ranks data according to a stereotypical form of alignment between the text and images, rather than some inherent measure of their quality. In this example, a common word paired with one gender is included by the filter more often than the same word paired with a different gender. Again, we see the feminization of the Asian identity (Azhar et al., 2021) and the diminution of women as “girls” (Madsen, 2021). We also observe that text containing the word “person” is more likely to pass the filter when paired with words related to men than when paired with words related to women. Other common words with higher man-related pass rates illustrate the ongoing stereotypes of associating men with technology and careers, which have been shown to be present among both humans and machine learning models (Oldenziel, 1999; Zuo and Tang, 2000; Caliskan et al., 2022).
@@ 5.2. Image group
Given the established biases and shortcomings of pre-trained third-party attribute predictors (Buolamwini and Gebru, 2018), in addition to using Amazon Rekognition, we introduce another technique to complement the previously presented results. We follow methodology very similar to Bianchi et al. (2023) and Luccioni et al. (2024), which uses embedding clustering techniques on a reference database (in our case, the Chicago Face Database with self-reported gender and race annotations (Ma et al., 2015)) in order to assess a model’s internal representation of gender and race. We defer to Section A.2 the additional details about the implementation and validation of this method in comparison to Amazon Rekognition. This method requires face box annotations, so we apply this method to the same set of $11,000$ images detected by Rekognition as containing a single face. Going forward, we refer to this technique as CLIP representation-based kNN, or CLIP kNN for short, and we follow the same annotation group names as used in the Chicago Face Database.
@@ 6.1. Language
For languages outside of English, Figure 7 shows the pass rates of each language (with at least $1,000$ samples) against their frequency in CommonPool. We observe that the next highest pass rates include Western languages like Spanish, Dutch, French, Catalan, German, and Portuguese, potentially due to their similarity to the English language. Furthermore, we demonstrate a similar notion of exclusion amplification as the trend from Section 5.2: data from detected languages that have lower representation in the raw dataset are discarded by the filter at higher rates (more details, along with a a more comprehensive set of languages, are presented in Section B.1).
@@ 6.3. IP address geolocation
In addition to country domain, we refer to an IP address geolocation database in order to obtain the country related to a source domain. Similar to the approach in Dodge et al. (2021), we use the IP2Location Lite IP-Country Database (IP2Location, 2024), which has been demonstrated to have high country-level accuracy (Komosny et al., 2017; Poese et al., 2011). We obtain the IP address set using the python socket library for a random selection of 1 million CommonPool samples, which result in 200 thousand unique websites. IP2Location then successfully identifies the country of 95.7% of the sampled website domains, which corresponds to 770 thousand CommonPool samples. We recognize that websites can be hosted in various data servers around the world, or based on user location, so IP addresses may not accurately represent the location of a website (Livadariu et al., 2020). As such, we use geolocation to complement the other methods we implement to obtain geographic region.
6.3.1. Most websites are hosted in the United States, and data hosted in Western countries bypass the filter at higher rates.
Figure 9 demonstrates that for both the raw dataset and filtered dataset, an overwhelming majority of data comes from the United States, although this could be due to the origin location used to ping the website. We also find that the countries with the higher pass rates correspond to those in Western regions, and that data hosted in non-Western regions are filtered out at higher rates.
Figure 9. Raw dataset frequency by country of IP address sorted in descending order by pass rate (included in parentheses). We display countries with at least 5,000 samples, and most of these non-Western countries (highlighted) have low pass rates.
@@
Bar graph of the frequency in the raw dataset by country of IP address. The USA has the highest frequency of around $450,000$, and all other countries are below $100,000$. Pass rates range from $0.43$ (India) to $0.10$ (Japan).

### v2

# {textblock*}

@@ 1. Introduction
Text and image datasets in machine learning have grown to the scale of billions (Devlin et al., 2018; Radford et al., 2019; Schuhmann et al., 2022) in order to build larger text and image generative models (Hoffmann et al., 2022; AI, 2024). To reach such magnitudes, datasets are being increasingly scraped from the web (Raffel et al., 2020; Gadre et al., 2024), through archives such as Common Crawl (Crawl, 2024). While data from these web dumps are expansive, they are not expertly curated and may reduce the efficacy of downstream models. Manually assessing every data point is infeasible, however, so this requires at-scale methods to remove data that curators deem undesirable.
To address these issues, machine learning practitioners and researchers have increasingly begun to rely on automated data filtering to improve training efficiency and discard what they refer to as “low-quality” data. Specifically, for multimodal data, researchers have applied a pretrained CLIP model to filter image-text data obtained from Common Crawl. As depicted in Figure 1, this CLIP-filtering method (referring to the use of the OpenAI CLIP model (Radford et al., 2021)) assesses similarity between an image and its corresponding text within the model’s embedding space. This filtering technique was used to create the LAION-400M (Schuhmann et al., 2021), LAION-5B (Schuhmann et al., 2022), and DataComp-1B (Gadre et al., 2024) datasets with sizes of 400 million, 5 billion, and 1 billion respectively, which were then used to train other open-source CLIP models. Because these models obtained state-of-the-art performance on several zero-shot image classification tasks, researchers have concluded CLIP-filtering to be the “most performant” method, as opposed to other filtering techniques that use caption length or image size (Gadre et al., 2024; Fang et al., 2023).
@@
At the same time, it is unclear what role CLIP-filtering plays on downstream text-to-image models or whether CLIP-filtering simply replicates the model’s training dataset, especially given the demographic biases embedded in OpenAI CLIP itself (Agarwal et al., 2021). While there have been speculations on the demographic biases of CLIP-filtering (Birhane et al., 2021), there has not been any exploration to assess exactly how the filter changes the demographic makeup of the training dataset. As such, we choose to examine this initial stage of dataset curation because of its potential impact on downstream models.
@@ 1.1. Contributions
We break down our investigation to address the following three questions: First: Which demographic groups are disproportionately excluded by CLIP-filtering? Different demographic groups being filtered at different rates implies a form of representational harm (Shelby et al., 2023). A filter that systematically omits data from a certain group may therefore worsen the performance of the resulting model on that group. As a result, we aim to quantify this breakdown between unfiltered and filtered data across various demographic dimensions. The need to assess sociodemographic attributes across multiple modalities motivates our next question: How can we measure sociodemographic attributes at scale within an image-text dataset? We examine a broad range of factors like gender, age, race, religion, sexuality, language, or geographic region. While by no means comprehensive, we infer attributes that appear in the text and image components that can easily be related to people. In addition to examining filtering discrepancies by demographics, we also examine discrepancies by data source: What types of websites are considered “high-quality” by the CLIP filter and what are the implications of certain websites being included? To the extent that it is valuable to know whether datasets are globally representative, we ask whether data from certain geographic regions or time periods of the internet are treated differently by CLIP-filtering, as well as what types of websites have data that are more likely to pass through the filter.
@@
    Filter discrepancies: Across multiple modalities, different imputed sociodemographic groups are filtered at different rates. For example, CLIP-filtering is more likely to exclude data relating to LGBTQ+ identities and non-Western regions.
@@
    Evaluation tools: To address our research questions, we also develop novel methodologies to effectively audit an image-text dataset at scale. We make our analysis code available at https://github.com/hongrachel/clip-filtering-bias/.
Overall, this existing data filtering method is a clear example of a value-embedded machine learning practice (Birhane et al., 2022). From our results, we see that CLIP-filtering does in fact substantially change the makeup of the final dataset. This reinforces more broadly the idea from Gururangan et al. (2022) that data filtering must encode a specific ideology of what constitutes as “high-quality data.” Prior works on image-text filtering assume that “quality” is often intrinsic to the data and that this “noise” to be weeded out is vaguely defined (Wenzek et al., 2019). The LAION and DataComp authors demonstrate the efficacy of the CLIP filter by selecting specific downstream metrics for specific datasets such as zero-shot ImageNet classification accuracy, rather than the societal impacts of downstream models (Schuhmann et al., 2021; Gadre et al., 2024). Here, we investigate a case where filtering is designed around particular assumptions of what data quality and performance should be without additional justification. Despite the supposed objectivity of these decisions, our findings prove that filtering embeds societal norms as to what should and should not be excluded. Following our analysis, we further elaborate on these implications and give recommendations for designing data filters in the future.
@@ 3. Related Work
In this section, we highlight previous work on bias audits and data filtering methods that are relevant to our study. We aim to contribute to the existing discussion on data filtering bias through a novel case study of data filtering on a popular large-scale dataset.
@@ 3.2. CLIP-filtered dataset audits
Some recent works have audited the LAION datasets in order to understand the nature of the content within these large-scale datasets. Recently, Thiel (2023) uses hash-based detection and other computational techniques to identify thousands of CSAM (Child Sexual Abuse Material) images in LAION-5B, which resulted in the takedown of the LAION-5B dataset in December, 2023 (Verma and Harwell, 2023). In addition, Birhane et al. (2021, 2023) have found significant rates of hateful content, explicit images, stereotypes, and racial slurs in both the LAION-400M and LAION-2B-en datasets. Given the biases of CLIP, they illustrate hypothetical examples in which misogynist or racist descriptions would pass the CLIP filter but reasonable, benign descriptions would be excluded (Birhane et al., 2021). The question then remains: What, exactly, does CLIP-filtering include and exclude? We extend their findings through examining the impact of filtering on demographic representation at scale.
@@ 3.3. Data filtering bias
There also has been much discussion on the systematic exclusion of data from marginalized identities in the data cleaning stage more broadly. Bender et al. (2021) caution early on that filtering can suppress discourse from marginalized identities and compound power imbalances in text data collection practices. From a sociotechnical lens, Muller and Strohmayer (2022) describe how data cleaning can be a method of erasure that is difficult to detect. Desai et al. (2023) draw on an archival perspective to examine the social value judgements embedded in filtering choices. We are inspired by prior filtering discussions and extend their approaches to a commonly used multimodal data pipeline. To our knowledge, there has been no work assessing data filtering bias at scale in the context of image-text datasets.
@@ 3.5. Downstream text-to-image model harms
There has been ample work that evaluate the societal issues of downstream models trained on CLIP-filtered data. Luccioni et al. (2024), for instance, find that Stable Diffusion (which trains on LAION-5B obtained via CLIP-filtering (Rombach et al., 2022)), underrepresents marginalized identities in their image generations. Moreover, researchers have demonstrated that Stable Diffusion generates unsafe, hateful, or sexualized content at significantly high rates (Qu et al., 2023; Wolfe et al., 2023). Exposure to stereotypical and problematic imagery in general has been shown to shape people’s beliefs and behavior (Branley and Covey, 2017; Slusher and Anderson, 1987), and this becomes more pressing as these models grow increasingly popular. As such, it thus becomes necessary to understand exactly where societal harms may arise in the ML pipeline.
@@ 4. Approach
We first replicate the CLIP-filtering step for the small version of CommonPool of 12.8 million image-text pairs. We use the CLIP ViT-L/14 model and set the threshold as $0.243$ to select the top $30\%$ of pairs according to CLIP similarity score as done in prior work (Schuhmann et al., 2022; Gadre et al., 2024), resulting in a filtered dataset with 3.84 million samples. In this section, we describe how we obtain demographic annotations and measure filtering discrepancies.
@@ 4.1. Imputed demographic group annotations
We determine demographic group or geographic region across multiple modalities in the CommonPool dataset: from the text content, the image content, and the source URL of the sample. As described in Section 5, this includes methods such as searching for mentions of keywords relating to demographic identity or applying an external face attribution predictor. Due to the sample size considered, we separate our analysis by modality, despite a sample containing both image and text, and leave the image-text intersection for follow-up work. We emphasize these findings do not represent behavior of filtering according to true demographics; instead, they reveal how filtering interacts with imputed demographic attributes. While we elaborate on the limitations of imputed demographics in Section 4.3, findings here still have broad implications, as imputed demographics can correlate, although imperfectly, with socially salient populations (Elliott et al., 2009).
@@ 4.3. Limitations
We recognize numerous limitations to the sociodemographic imputation and source analysis techniques that we use. Third-party gender predictors, for instance, have been demonstrated to have disparate error rates across sensitive attributes and can rely on biases present in the face detection step (Buolamwini and Gebru, 2018; Schwemmer et al., 2020). Furthermore, these predictors only give binary labels as “Male” and “Female” which conflates gender with sex and ignores non-binary individuals (Valdes, 1996). Identity keywords in the text also are rough proxies for the presence of sociodemographic attributes in the image or text. We acknowledge that gender and race are not inherently visual components, but rather socioculturally defined (Butler, 2013; Lopez, 1995).
@@ 5. Demographic Group Analysis
This section describes the demographic group imputation methods along both text and image modalities and presents their corresponding results. We include additional analysis in Appendix B.
@@ 5.1. Identity keyword
(a) Pass rate by identity keyword sorted in descending order.
@@
The aforementioned gender stereotype associations provide evidence that CLIP-filtering ranks data according to a stereotypical form of alignment between the text and images, rather than some inherent measure of their quality. In this example, a common word paired with one gender is included by the filter more often than the same word paired with a different gender. We see the feminization of the Asian identity (Azhar et al., 2021) and the diminution of women as “girls” (Madsen, 2021). We also observe that text containing the word “person” is more likely to pass the filter when paired with words related to men than when paired with words related to women. Other common words with higher man-related pass rates illustrate the ongoing stereotypes of associating men with technology and careers, which have been shown to be present among both humans and machine learning models (Oldenziel, 1999; Zuo and Tang, 2000; Caliskan et al., 2022).
@@ 5.2. Image group
Given the established biases and shortcomings of pre-trained third-party attribute predictors (Buolamwini and Gebru, 2018), in addition to using Amazon Rekognition, we introduce another technique to complement the previously presented results. We follow methodology very similar to Bianchi et al. (2023) and Luccioni et al. (2024), which uses embedding clustering techniques on a reference database (in our case, the Chicago Face Database using their self-reported gender and race annotations (Ma et al., 2015)) in order to assess a model’s internal representation of gender and race. We defer to Section A.2 the additional details about the implementation and validation of this method in comparison to Amazon Rekognition. This method requires face box annotations, so we apply this method to the same set of $11,000$ images detected by Rekognition as containing a single face. Going forward, we refer to this technique as CLIP representation-based kNN, or CLIP kNN for short, and we follow the same annotation group names as used in the Chicago Face Database.
@@ 6.1. Language
For languages outside of English, Figure 7 shows the pass rates of each language (with at least $1,000$ samples) against their frequency in CommonPool. We observe that the next highest pass rates include Western languages like Spanish, Dutch, French, Catalan, German, and Portuguese, potentially due to their similarity to the English language. Furthermore, we demonstrate a similar notion of exclusion amplification as the trend from Section 5.2: data from detected languages that have lower representation in the raw dataset are discarded by the filter at higher rates (more details, along with a a more comprehensive set of languages, are presented in Section B.1.
@@ 6.3. IP address geolocation
In addition to country domain, we refer to an IP address geolocation database in order to determine the country related to a source domain. Similar to the approach in Dodge et al. (2021), we use the IP2Location Lite IP-Country Database (IP2Location, 2024), which has been demonstrated to have high country-level accuracy (Komosny et al., 2017; Poese et al., 2011). We obtain the IP address set using the python socket library for a random selection of 1 million CommonPool samples, which results in 200 thousand unique websites. IP2Location then successfully identifies the country of 95.7% of the sampled website domains, which corresponds to 770 thousand CommonPool samples. We recognize that websites can be hosted in various data servers around the world, or based on user location, so IP addresses may not accurately represent the location of a website (Livadariu et al., 2020). As such, we use inferred geolocation to complement the other methods we implement to obtain geographic region.
6.3.1. The inferred geolocation of most websites come from the United States, and samples with IP address locations from Western countries bypass the filter at higher rates.
Figure 9 demonstrates that for both the raw dataset and filtered dataset, the inferred geolocation of an overwhelming majority of data come from the United States, although this may be due to the origin location used to ping the website. We also find that the IP address locations of countries with higher pass rates correspond to those in Western regions, and that data with inferred geolocations from non-Western regions are filtered out at higher rates.
Figure 9. Raw dataset frequency by country of IP-based location sorted in descending order by pass rate (included in parentheses). We display countries with at least 5,000 samples, and most of these non-Western countries (highlighted) have low pass rates.
@@
Bar graph of the frequency in the raw dataset by country of inferred geolocation. The USA has the highest frequency of around $450,000$, and all other countries are below $100,000$. Pass rates range from $0.43$ (India) to $0.10$ (Japan).